The resultant of two forces one double the other in magnitudes is perpendicular to the smaller of the two forces.The angle between the two forces is

Grade:10

2 Answers

Arun
4 years ago
Let two forces F₁ and F₂ are given
A/C to question, 
|F₁| = 2|F₂|
Resultant of two forces F₁ and F₂ are given by 
R = F₁ + F₂ 
A/C to question, 
R is perpendicular upon smaller force.
Here smaller force is F₂.
So, R.F₂ = 0
⇒(F₁ + F₂).F₂ = 0
⇒F₁.F₂ + F₂² = 0 
⇒|F₁||F₂|cosФ + |F₂|² = 0 
Now, put |F₁| = 2|F₂| 
⇒2|F₂|.|F₂|cosФ + |F₂|² = 0
⇒2cosФ + 1 = 0 
⇒cosФ = -1/2 = cos120°
Hence, angle between two forces is 120°
 
Khimraj
4 years ago
Assume that F and 2F respectively
Let R be the Resultant .
It is given that R is perpendicular to F
Then
R^2+ F^2 = 4F^2
R^2 = 3F^2
3F^2-5F^2 = 4F^2cosα
-2F^2 = 4F^2cosα
Cosα = -1/2
α = 2π/3.